# Rings of non-commutative functions and their fields of fractions

### v1

@@ 6.1 Strong convergence of meromorphic functions
Let us fix a free semicircular tuple $s_{1},\ldots,s_{d}$. Let $\operatorname{SGRM}(n,\tfrac{1}{n})$ denote the class of all $n\times n$ self-adjoint random matrices on a probability space $\Omega$ with independent and identically distributed (i.i.d.) Gaussian entries with mean 0 and variance $\frac{1}{n}$. The normalization of the variance goes back to Voiculescu [51]. In [20, Theorem A] Haagerup and Thorbjørnsen proved that if we take a sequence $(X_{1}^{(n)},\cdots,X_{d}^{(n)})$ of $d$-tuples of independent elements of $\operatorname{SGRM}(n,\tfrac{1}{n})$ for all $n\in\mathbb{N}$, then there exists a measure zero set $\Xi\subset\Omega$, such that for every $p\in\mathbb{C}\langle z\rangle$ and every $\omega\in\Omega\setminus\Xi$,
$\lim_{n\to\infty}\left\|p(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega)\right\|=\|p(s_{1},\cdots,s_{d})\|.$
This result was extended to matrix-valued polynomials by Haagerup, Schultz, and Thorbjørnsen in [19]. Using the terminology of [54], we will say that the sequence $(X_{1}^{(n)},\cdots,X_{d}^{(n)})$ converges strongly to the free semicircular tuple $s_{1},\cdots,s_{n}$. In [54, Corollary 1], Yin showed that if $\mathfrak{r}$ is a rational expression with the tuple $(s_{1},\cdots,s_{d})$ in its domain, then for $n$ big enough and almost every $\omega\in\Omega$, $(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega)$ lies in the domain of $\mathfrak{r}$ and
$\lim_{n\to\infty}\left\|\mathfrak{r}(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|=\|\mathfrak{r}(s_{1},\ldots,s_{d})\|.$
@@
$\lim_{n\to\infty}\mathbb{E}\left(\frac{1}{n}\operatorname{tr}(\mathfrak{r}(X_{1}^{(n)},\ldots,X_{d}^{(n)})\right)=\tau(\mathfrak{r}(s_{1},\ldots,s_{n})).$
Here $\mathbb{E}$ stands for the expectation of a random variable. This result extends the result of Haagerup and Thorbjørnsen to the free skew field. Our goal is to extend these results even further to the universal skew field of fractions, $\mathscr{M}_{d}$, of $\mathscr{O}_{d}$. Note that by the result of Haagerup and Thorbjørnsen, the row tuple, $(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))$ is, almost surely, in a row ball of bounded radius. We start with a simple lemma.
@@
$\lim_{n\to\infty}\left\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega)\right\|=\|f(s_{1},\cdots,s_{d})\|.$
@@
Any such $f$ is given by a free FPS expansion of $f$ about 0,
$f(\mathfrak{z})=\sum_{k=0}^{\infty}\sum_{|\alpha|=k}\hat{f}_{\alpha}\mathfrak{z}^{\alpha}.$
Let $p_{m}=\sum_{k=0}^{m}\sum_{|\alpha|=k}\hat{f}_{\alpha}\mathfrak{z}^{\alpha}$. Then, this sequence of free polynomial partial sums, $p_{m}$, converges uniformly to $f$ on any NC row-ball of finite radii. Now by [20, Theorem A], for almost all $\omega\in\Omega$ and all $m\in\mathbb{N}$, we have that
@@
$\left|\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\|-\|f(s_{1},\cdots,s_{d})\|\right|\leq\left\|f(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))-p_{m}(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|$
@@
For every $k\in\mathbb{N}$ and every full matrix $F\in\mathscr{O}_{d}^{k\times k}$, if $F(s_{1},\cdots,s_{d})$ is invertible, then for almost every $\omega\in\Omega$ for $n$ sufficiently large, $F(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))$ is invertible.
@@
We note that there is a natural anti-holomorphic involution on the free algebra $\mathbb{C}\langle\mathfrak{z}\rangle$, such that the variables are self-adjoint (namely, $\mathfrak{z}_{j}^{*}=\mathfrak{z}_{j}$ for all $1\leq j\leq d$.) The involution is given by transposing the monomials and taking the complex conjugate of the coefficients. This involution extends naturally to $\mathscr{O}_{d}$ and to matrices over $\mathscr{O}_{d}$. Let $F\in\mathscr{O}_{d}^{k\times k}$, then $F^{*}F\in\mathscr{O}_{d}^{k\times k}$ is self-adjoint. Moreover, for every tuple of self-adjoint elements $Y_{1},\cdots,Y_{d}\in B(\mathcal{H})$, we have that $(F^{*}F)(Y_{1},\ldots Y_{d})=F(Y_{1},\ldots,Y_{d})^{*}F(Y_{1},\cdots,Y_{d})$. To simplify notations, let $T=F(s_{1},\cdots,s_{d})$ and $T_{n}(\omega)=F(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))$. We note that a self-adjoint operator $A\in B(\mathcal{H})$ is invertible if and only if $A^{*}A$ is invertible if and only if $\big\lVert\lVert A^{2}\rVert I-A^{*}A\big\rVert<\|A\|^{2}$. By the previous lemma $\|T_{n}(\omega)\|\to\|T\|$ almost surely. Moreover, almost surely
$\lim_{n\to\infty}\Big\lVert\lVert T\rVert^{2}I-T_{n}(\omega)^{*}T_{n}(\omega)\Big\rVert=\Big\lVert\lVert T\rVert^{2}I-T^{*}T\Big\rVert$
Now
@@
The following lemma is an immediate consequence of the fact that the group of invertible operators on a complex Hilbert space is operator-norm open and the inversion map is uniformly continuous on this group.
@@
Let $F\in\mathscr{O}_{d}^{k\times k}$ be a full matrix. Let $Q_{m}\in\mathbb{C}\langle\mathfrak{z}\rangle^{k\times k}$ be the partial sums of the FPS expansion of $F$ around the origin. Then, for every $X=(X_{1},\cdots,X_{d})\in\mathscr{B}(\mathcal{H})^{d}$, such that $F(X)$ is invertible, $Q_{m}(X)$ is invertible for $m$ big enough. Moreover, for every such $X$, there exists $r>0$, such that $Q_{m}^{-1}\to F^{-1}$ uniformly on the closed uniform NC row-ball of radius $r$ centered at $X$.
It follows immediately that if $f=d-b^{t}A^{-1}c$ is a rational expression in $\mathscr{M}_{d}$, then there exists a sequence of rational expressions $\mathfrak{r}_{m}$, such that $\mathfrak{r}_{m}\to f$ uniformly on uniform balls centered at points in the domain of the rational expression. Indeed, we can approximate $b$, $c$, and $d$ uniformly on subballs by polynomial vectors and the previous lemma takes care of the rest.
@@
$\lim_{n\to\infty}\left\|f(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))\right\|=\|f(s_{1},\ldots,s_{d})\|.$
@@
By Theorem 6.1 and its preceding discussion, we can choose a representation of $f=b^{*}A^{-1}c$, such that $b,c\in\mathbb{C}^{k}$, $A\in\mathscr{O}_{d}^{k\times k}$ is full, and $A(s_{1},\ldots,s_{d})$ is invertible. By Lemma 6.5, $A(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ is invertible, for almost every $\omega$ and $n$ big enough. Now fix $\omega\in\Omega\setminus\Xi$, where $\Xi$ is the measure 0 set that we must remove to ensure convergence. By the observation preceding the statement of the theorem, there exists a sequence of rationals $\mathfrak{r}_{m}$ that converges to $f$ uniformly on uniform balls centered at points in the domain of $f$. It remains to perform a calculation. For $m$ big enough, the following inequality makes sense
@@
Given $\varepsilon>0$, we can choose $m$ big enough, so that
@@
$\limsup_{n\to\infty}\left\lvert\lVert f(s_{1},\ldots,s_{d})\rVert-\lVert f(X_{1}^{(n)}(\omega),\ldots,X_{d}^{(n)}(\omega))\rVert\right\rvert\leq\varepsilon.$
@@
$\lim_{n\to\infty}\mathbb{E}\left(\frac{1}{n}\operatorname{tr}\,f(X_{1}^{(n)},\cdots,X_{d}^{(n)})\right)=\tau\left(f(s_{1},\ldots,s_{n})\right).$
@@ 7 Outlook
We leave several interesting questions open. One natural question arises from the fact that the Fréchet topology on $\mathscr{O}_{d}$, viewed as an inverse limit of NC disk algebras, is independent of the chosen operator space structure on $\mathbb{C}^{d}$. Here, we have chosen to work with the row operator space structure over $\mathbb{C}^{d}$ to exploit non-commutative Hardy space techniques, since the free Hardy space can be viewed as a Hilbert space of uniformly analytic NC functions on the NC unit row-ball centred at $0\in\mathbb{C}^{1\times d}$. However, our methods depend crucially on the ‘row operator space structure’ that is implicit in the definition of Popescu’s NC disk algebra as NC functions in the NC unit row-ball. (The column operator space structure for $\mathbb{C}^{d}$ would work equally well.) If $\mathcal{E}$ is an operator space structure on $\mathbb{C}^{d}$, then we denote by $\mathbb{B}_{\mathcal{E}}^{d}$ the NC set that is the open unit ball of $\mathcal{E}$. The algebras $\mathbb{H}^{\infty}(\mathbb{B}_{\mathcal{E}}^{d})$ and $\mathbb{A}_{d}(\mathcal{E})$ were studied by Sampat and Shalit [44, 45]. One can ask whether the algebras $\mathscr{O}(R\cdot\mathbb{B}_{\mathcal{E}}^{d})$ given by NC functions that are uniformly bounded on $r\cdot\overline{\mathbb{B}_{\mathcal{E}}^{d}}$ for every $0<r<R$ are semifirs that satisfy an analogue of the analytic Bergman Nullstelensatz.
A second natural question is whether every closed right (respectively, left) ideal in $\mathscr{O}_{d}$ is free as a right (respectively, left) $\mathscr{O}_{d}$-module in the sense of completed tensor products as in [47].
@@
In [6], the authors have introduced a second skew field of NC meromorphic functions, $\mathscr{M}_{d}^{\mathscr{C}}$. This is the universal skew field of fractions of the semifir, $\mathscr{O}_{d}^{\mathscr{C}}$ of all NC functions that admit a compact realization, $(A,b,c)\in\mathscr{C}(\mathcal{H})^{d}\times\mathcal{H}\times\mathcal{H}$, where $\mathcal{H}$ is a complex, separable Hilbert space and $\mathscr{C}(\mathcal{H})$ denotes the compact linear operators on $\mathcal{H}$. That is, given any $h\in\mathscr{O}_{d}^{\mathscr{C}}$, given by the compact realization $(A,b,c)$, and any $Z\in\mathbb{C}^{(n\times n)\cdot d}$ for which the monic, affine linear pencil, $L_{A}(Z):=I_{n}\otimes I_{\mathcal{H}}-\sum_{j=1}^{d}Z_{j}\otimes A_{j}$, is invertible,
@@
By [6, Theorem 4.5 and Corollary 4.4], any free formal power series, $f\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$, is a uniformly entire NC function in $\mathscr{O}_{d}$ if and only if it has a compact and jointly quasinilpotent realization, $(A,b,c)$. That is, each component operator, $A_{j}\in\mathscr{C}(\mathcal{H})$ is compact and the completely positive map, $\mathrm{Ad}_{A,A^{*}}:\mathscr{B}(\mathcal{H})\rightarrow\mathscr{B}(\mathcal{H})$, of adjunction by the $A_{j}$ and $A_{j}^{*}$, is quasinilpotent as an element of the unital Banach algebra, $\mathscr{B}(X)$, where $X:=\mathscr{B}(\mathcal{H})$. In particular, $\mathscr{O}_{d}\subsetneqq\mathscr{O}_{d}^{\mathscr{C}}$. It then follows from the standard realization algorithm for sums, products and inverses of NC functions with realizations that a univariate formal power series is the Taylor series of a globally meromorphic function (which is analytic at 0) if and only if it has a compact realization [6, Theorem 5.1 and Theorem 5.3]. This motivates the interpretation of $\mathscr{O}_{d}^{\mathscr{C}}$, the semifir of all NC functions with compact realizations, as the ring of globally meromorphic NC functions with analytic germs at 0, as well as the interpretation of its universal skew field, $\mathscr{M}_{d}^{\mathscr{C}}$ as the skew field of globally meromorphic NC functions. Here, elements of $\mathscr{M}_{d}^{\mathscr{C}}$ can indeed be identified with bona fide uniformly analytic NC functions whose domains are uniformly open subsets that are level-wise analytic-Zariski dense at all sufficiently high levels in the NC universe, see [6, Section 6.1 and Theorem 6.4].
In one-variable, it is not difficult to see, from the results of [6], that these two definitions of the field of globally meromorphic functions are the same, i.e. that $\mathscr{M}_{1}=\mathscr{M}_{1}^{\mathscr{C}}$. Therefore, it would be very interesting to know whether these two skew fields of “NC meromorphic functions”, $\mathscr{M}_{d}$ and $\mathscr{M}_{d}^{\mathscr{C}}$ are actually the same field for any $d\in\mathbb{N}$. Since, as discussed above, any $h\in\mathscr{O}_{d}$ has a compact and jointly quasinilpotent realization, we have an embedding $\mathscr{O}_{d}\hookrightarrow\mathscr{O}_{d}^{\mathscr{C}}$. If it could be proven that this embedding, or the embedding $\mathscr{O}_{d}\hookrightarrow\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ is honest, it would follow that $\mathscr{M}_{d}\hookrightarrow\mathscr{M}_{d}^{\mathscr{C}}$, so that $\mathscr{M}_{d}$ could be identified with a skew sub-field of $\mathscr{M}_{d}^{\mathscr{C}}$. A special case of a classical question posed by Henri Poincaré, and ultimately answered, in the affirmative by the positive solution to the Cousin problems, is whether any globally meromorphic function in $\mathbb{C}^{d}$ is a ratio of entire functions [31]. Thus, we arrive at a “NC Poincaré problem:”

### v2

@@ 6.1 Strong convergence of meromorphic functions
Let us fix a free semicircular tuple $s_{1},\cdots,s_{d}$. Let $\operatorname{SGRM}(n,\tfrac{1}{n})$ denote the class of all $n\times n$ self-adjoint random matrices on a probability space $\Omega$ with independent and identically distributed (i.i.d.) Gaussian entries with mean 0 and variance $\frac{1}{n}$. This normalization of the variance goes back to Voiculescu [53]. In [21, Theorem A] Haagerup and Thorbjørnsen proved that if we take a sequence $(X_{1}^{(n)},\cdots,X_{d}^{(n)})$ of $d$-tuples of independent elements of $\operatorname{SGRM}(n,\tfrac{1}{n})$ for all $n\in\mathbb{N}$, then there exists a measure zero set $\Xi\subset\Omega$, such that for every $p\in\mathbb{C}\langle z\rangle$ and every $\omega\in\Omega\setminus\Xi$,
$\lim_{n\to\infty}\left\|p(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|=\|p(s_{1},\cdots,s_{d})\|.$
This result was extended to matrix-valued polynomials by Haagerup, Schultz, and Thorbjørnsen in [20]. Using the terminology of [56], we will say that the sequence $(X_{1}^{(n)},\cdots,X_{d}^{(n)})$ converges strongly to the free semicircular tuple $s_{1},\cdots,s_{n}$. In [56, Corollary 1], Yin showed that if $\mathfrak{r}$ is a rational expression with the tuple $(s_{1},\cdots,s_{d})$ in its domain, then for $n$ big enough and almost every $\omega\in\Omega$, $(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ lies in the domain of $\mathfrak{r}$ and
$\lim_{n\to\infty}\left\|\mathfrak{r}(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|=\|\mathfrak{r}(s_{1},\cdots,s_{d})\|.$
@@
$\lim_{n\to\infty}\mathbb{E}\left(\frac{1}{n}\operatorname{tr}(\mathfrak{r}(X_{1}^{(n)},\cdots,X_{d}^{(n)}))\right)=\tau(\mathfrak{r}(s_{1},\cdots,s_{n})).$
Here, $\mathbb{E}$ stands for the expectation of a random variable. This result extends the result of Haagerup and Thorbjørnsen to the free skew field. Our goal is to extend these results even further to the universal skew field of fractions, $\mathscr{M}_{d}$, of $\mathscr{O}_{d}$. Note that by the result of Haagerup and Thorbjørnsen, the row tuple, $(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ is, almost surely, in a row ball of bounded radius. We start with a simple lemma.
@@
$\lim_{n\to\infty}\left\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|=\|f(s_{1},\cdots,s_{d})\|.$
@@
Any such $f$ is given by a free FPS,
$f(\mathfrak{z})=\sum_{k=0}^{\infty}\sum_{|\alpha|=k}\hat{f}_{\alpha}\,\mathfrak{z}^{\alpha},$
with infinite radius of convergence. Let $p_{m}=\sum_{k=0}^{m}\sum_{|\alpha|=k}\hat{f}_{\alpha}\mathfrak{z}^{\alpha}$ be the $m$-th free polynomial partial sum of $f$. Then, this sequence of free polynomial partial sums, $p_{m}$, converges uniformly to $f$ on any NC row-ball of finite radii. By [21, Theorem A], for almost all $\omega\in\Omega$ and all $m\in\mathbb{N}$,
@@
$\left|\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\|-\|f(s_{1},\cdots,s_{d})\|\right|\leq\left\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))-p_{m}(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|$
@@
For every $k\in\mathbb{N}$ and every full matrix $F\in\mathscr{O}_{d}^{k\times k}$, if $F(s_{1},\cdots,s_{d})$ is invertible, then for almost every $\omega\in\Omega$ for $n$ sufficiently large, $F(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ is invertible.
@@
There is a natural anti-holomorphic involution on the free algebra $\mathbb{C}\langle\mathfrak{z}\rangle$, with respect to which the formal NC variables are self-adjoint (namely, $\mathfrak{z}_{j}^{*}=\mathfrak{z}_{j}$ for all $1\leq j\leq d$). The involution is given by transposing the monomials and taking the complex conjugate of the coefficients,
$\sum_{\omega\in\mathbb{W}_{d}}\hat{p}_{\omega}\,\mathfrak{z}^{\omega}\qquad\mapsto\qquad\sum\overline{\hat{p}_{\omega}}\,\mathfrak{z}^{\omega^{\mathrm{t}}}.$
This involution extends naturally to $\mathscr{O}_{d}$ and to matrices over $\mathscr{O}_{d}$. Let $F\in\mathscr{O}_{d}^{k\times k}$, then $F^{*}F\in\mathscr{O}_{d}^{k\times k}$ is self-adjoint. Moreover, for every tuple of self-adjoint elements $Y_{1},\cdots,Y_{d}\in B(\mathcal{H})$, we have that $(F^{*}F)(Y_{1},\cdots,Y_{d})=F(Y_{1},\cdots,Y_{d})^{*}F(Y_{1},\cdots,Y_{d})$. To simplify notations, let $T=F(s_{1},\cdots,s_{d})$ and $T_{n}(\omega)=F(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$. We note that a self-adjoint operator $A\in B(\mathcal{H})$ is invertible if and only if $A^{*}A$ is invertible if and only if $\big\lVert\lVert A^{2}\rVert I-A^{*}A\big\rVert<\|A\|^{2}$. By the previous lemma $\|T_{n}(\omega)\|\to\|T\|$ almost surely. Moreover, almost surely
$\lim_{n\to\infty}\Big\lVert\lVert T\rVert^{2}I-T_{n}(\omega)^{*}T_{n}(\omega)\Big\rVert=\Big\lVert\lVert T\rVert^{2}I-T^{*}T\Big\rVert,$
and
@@
The following lemma is an immediate consequence of the fact that the group of invertible operators on a complex Hilbert space is operator–norm open and the inversion map is uniformly continuous on this group.
@@
Let $F\in\mathscr{O}_{d}^{k\times k}$ be a full matrix. Let $Q_{m}\in\mathbb{C}\langle\mathfrak{z}\rangle^{k\times k}$ be the partial sums of the free FPS expansion of $F$ around the origin. Then for every $X=(X_{1},\cdots,X_{d})\in\mathscr{B}(\mathcal{H})^{d}$ so that $F(X)$ is invertible, $Q_{m}(X)$ is invertible for $m$ big enough. Moreover, for every such $X$, there exists $r>0$, such that $Q_{m}^{-1}\to F^{-1}$ uniformly on the closed uniform NC row-ball of radius $r$ centered at $X$.
It follows immediately that if $f$ is an NC rational expression in $\mathscr{M}_{d}$, then there exists a sequence of rational expressions $\mathfrak{r}_{m}$, such that $\mathfrak{r}_{m}\to f$ uniformly on uniform balls centered at points in the domain of the rational expression. Indeed, we can approximate $b$, $c$, and $d$ uniformly on subballs by polynomial vectors and the previous lemma takes care of the rest.
@@
$\lim_{n\to\infty}\left\|f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\right\|=\|f(s_{1},\cdots,s_{d})\|.$
@@
By Theorem 6.1 and its preceding discussion, we can choose a representation of $f=b^{*}A^{-1}c$, such that $b,c\in\mathbb{C}^{k}$, $A\in\mathscr{O}_{d}^{k\times k}$ is full, and $A(s_{1},\cdots,s_{d})$ is invertible. By Lemma 6.6, $A(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))$ is invertible, for almost every $\omega$ and $n$ big enough. Now fix $\omega\in\Omega\setminus\Xi$, where $\Xi$ is the measure 0 set that we must remove to ensure convergence. By the observation preceding the statement of the theorem, there exists a sequence of rationals $\mathfrak{r}_{m}$ that converges to $f$ uniformly on uniform balls centered at points in the domain of $f$. It remains to perform a calculation. For $m$ big enough, the following inequality makes sense
@@
Given $\varepsilon>0$, we can choose $m$ big enough so that
@@
$\limsup_{n\to\infty}\left\lvert\lVert f(s_{1},\cdots,s_{d})\rVert-\lVert f(X_{1}^{(n)}(\omega),\cdots,X_{d}^{(n)}(\omega))\rVert\right\rvert\leq\varepsilon.$
@@
$\lim_{n\to\infty}\mathbb{E}\left(\frac{1}{n}\operatorname{tr}\,f(X_{1}^{(n)},\cdots,X_{d}^{(n)})\right)=\tau\left(f(s_{1},\cdots,s_{n})\right).$
@@ 7 Outlook
We leave several interesting questions open. One natural question arises from the fact that the Fréchet topology on $\mathscr{O}_{d}$, viewed as an inverse limit of NC disk algebras, is independent of the chosen operator space structure on $\mathbb{C}^{d}$. Here, we have chosen to work with the row operator space structure over $\mathbb{C}^{d}$ to exploit non-commutative Hardy space techniques, since the free Hardy space can be viewed as a Hilbert space of uniformly analytic NC functions on the NC unit row-ball centred at $0\in\mathbb{C}^{1\times d}$. However, our methods depend crucially on the ‘row operator space structure’ that is implicit in the definition of Popescu’s NC disk algebra as NC functions in the NC unit row-ball. (The column operator space structure for $\mathbb{C}^{d}$ would work equally well.) If $\mathcal{E}$ is an operator space structure on $\mathbb{C}^{d}$, then we denote by $\mathbb{B}_{\mathcal{E}}^{d}$ the NC set that is the open unit ball of $\mathcal{E}$. The algebras $\mathbb{H}^{\infty}(\mathbb{B}_{\mathcal{E}}^{d})$ and $\mathbb{A}_{d}(\mathcal{E})$ were studied by Sampat and Shalit [46, 47]. One can ask whether the algebras $\mathscr{O}(R\cdot\mathbb{B}_{\mathcal{E}}^{d})$ given by NC functions that are uniformly bounded on $r\cdot\overline{\mathbb{B}_{\mathcal{E}}^{d}}$ for every $0<r<R$ are semifirs that satisfy an analogue of the analytic Bergman Nullstellensatz. A second natural question is whether every closed right (respectively, left) ideal in $\mathscr{O}_{d}$ is free as a right (respectively, left) $\mathscr{O}_{d}$-module in the sense of completed tensor products as in [49].
@@
In [6], the authors have introduced a second skew field of NC meromorphic functions, $\mathscr{M}_{d}^{\mathscr{C}}$. This is the universal skew field of fractions of the semifir, $\mathscr{O}_{d}^{\mathscr{C}}$, of all NC functions that admit a compact realization, $(A,b,c)\in\mathscr{C}(\mathcal{H})^{d}\times\mathcal{H}\times\mathcal{H}$, where $\mathcal{H}$ is a complex, separable Hilbert space and $\mathscr{C}(\mathcal{H})$ denotes the compact linear operators on $\mathcal{H}$. That is, given any $h\in\mathscr{O}_{d}^{\mathscr{C}}$, given by the compact realization $(A,b,c)$, and any $Z\in\mathbb{C}^{(n\times n)\cdot d}$ for which the monic, affine–linear pencil, $L_{A}(Z):=I_{n}\otimes I_{\mathcal{H}}-\sum_{j=1}^{d}Z_{j}\otimes A_{j}$, is invertible,
@@
By [6, Theorem 4.5 and Corollary 4.4], any free formal power series, $f\in\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle$, is a uniformly entire NC function in $\mathscr{O}_{d}$ if and only if it has a compact and jointly quasinilpotent realization, $(A,b,c)$. That is, each component operator, $A_{j}\in\mathscr{C}(\mathcal{H})$ is compact and the completely positive map, $\mathrm{Ad}_{A,A^{*}}:\mathscr{B}(\mathcal{H})\rightarrow\mathscr{B}(\mathcal{H})$, of adjunction by the $A_{j}$ and $A_{j}^{*}$, is quasinilpotent as an element of the unital Banach algebra, $\mathscr{B}(X)$, where $X:=\mathscr{B}(\mathcal{H})$. In particular, $\mathscr{O}_{d}\subsetneqq\mathscr{O}_{d}^{\mathscr{C}}$. It then follows from the standard realization algorithm for sums, products and inverses of NC functions with realizations that a univariate formal power series is the Taylor series of a globally meromorphic function (which is analytic at 0) if and only if it has a compact realization [6, Theorem 5.1 and Theorem 5.3]. This motivates the interpretation of $\mathscr{O}_{d}^{\mathscr{C}}$, the semifir of all NC functions with compact realizations, as the ring of globally meromorphic NC functions with analytic germs at 0, as well as the interpretation of its universal skew field, $\mathscr{M}_{d}^{\mathscr{C}}$ as the skew field of globally meromorphic NC functions. Here, elements of $\mathscr{M}_{d}^{\mathscr{C}}$ can indeed be identified with bona fide uniformly analytic NC functions whose domains are uniformly open subsets that are level-wise analytic-Zariski dense at all sufficiently high levels in the NC universe, see [6, Section 6.1 and Theorem 6.4].
In one-variable, it is not difficult to see, from the results of [6], that these two definitions of the field of globally meromorphic functions are the same, i.e. that $\mathscr{M}_{1}=\mathscr{M}_{1}^{\mathscr{C}}$. Therefore, it would be very interesting to know whether these two skew fields of “NC meromorphic functions”, $\mathscr{M}_{d}$ and $\mathscr{M}_{d}^{\mathscr{C}}$ are actually the same field for any $d\in\mathbb{N}$. Since, as discussed above, any $h\in\mathscr{O}_{d}$ has a compact and jointly quasinilpotent realization, we have an embedding $\mathscr{O}_{d}\hookrightarrow\mathscr{O}_{d}^{\mathscr{C}}$. If it could be proven that this embedding, or the embedding $\mathscr{O}_{d}\hookrightarrow\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ is honest, it would follow that $\mathscr{M}_{d}\hookrightarrow\mathscr{M}_{d}^{\mathscr{C}}$, so that $\mathscr{M}_{d}$ could be identified with a skew sub-field of $\mathscr{M}_{d}^{\mathscr{C}}$. A special case of a classical question posed by Henri Poincaré, and ultimately answered, in the affirmative by the positive solution to the Cousin problems, is whether any globally meromorphic function in $\mathbb{C}^{d}$ is a ratio of entire functions [32]. Thus, we arrive at an “NC Poincaré problem”: